Two spring with the same force constant of 75 N/m are connected end to end, and the

combination is pulled out by a force of 20 N. a) What is the PE of the stretched springs?

b) If the two springs are connected side-by-side and pulled by the same force, what will be it’s

PE now?


1
Expert's answer
2021-05-25T14:25:14-0400

By definition, PE of the stretched springs is given as follows:


"PE = \dfrac{kx^2}{2}"

where "k" is the total force constant of two springs, and "x" is their elongation.

a) In case of series connection the total spring constant will be the following (see https://socratic.org/questions/what-is-the-spring-constant-in-parallel-connection-and-series-connection):


"k = \dfrac{k_1 k_2}{k_1+k_2} = \dfrac{k_1}{2}"

where "k_1 = k_2 = 70N\/m" are the spring constants of two springs.

The elongation is given by the Hook's law:


"x = \dfrac{F}{k} = \dfrac{2F}{k_1}"

where "F = 20N" is the applied force. Thus, obtain:


"PE = \dfrac{k_1\cdot (2F/k_1)^2}{2\cdot 2} = \dfrac{F^2}{k_1}\\
PE = \dfrac{20^2}{70} \approx 5.7J"

b) For the springs in parallel, have:


"k = k_1 + k_2 = 2k_1\\
x = \dfrac{F}{k} = \dfrac{F}{2k_1}\\
PE = \dfrac{2k_1\cdot (F/2k_1)^2}{2} = \dfrac{F^2}{4k_1}\\
PE = \dfrac{20^2}{4\cdot 70} \approx 1.4J"

Answer. a) 5.7 J, b) 1.4 J.